# Relativistic framework for high-precision GNSS-based navigation in cislunar space

### v1

@@ I Introduction
For interplanetary and lunar trajectories, spacecraft dynamics are governed by the Einstein–Infeld–Hoffmann equations in the BCRS [7, 1]. The differential rate between Terrestrial Time (TT) and Barycentric Coordinate Time (TCB), $1.5505\times 10^{-8}\,\mathrm{s/s}$, yields $\sim 1.34\,\mathrm{ms/day}$ offsets near Earth, significantly affecting one‐ and two‐way light‐time calculations [8, 9]. In strong‐field regions—e.g. at ten solar radii where the Parker Solar Probe operates—the relativistic correction to Newtonian acceleration ($\sim 1.22\times 10^{-6}\,\mathrm{m/s^{2}}$) becomes non‐negligible relative to the $2.74\,\mathrm{m/s^{2}}$ background [10, 11]. Even at 1 AU, frame‐dragging and potential‐induced time dilation ($\sim 9.87\times 10^{-9}\,\mathrm{s/s}$) must be applied when transforming between geocentric and barycentric frames to maintain picosecond‐level timing consistency.
As GNSS technology advances with the goal of achieving cm-level positional accuracy, the inclusion of general relativistic corrections in orbit determination models becomes indispensable [12, 13, 14, 15, 16]. These corrections affect multiple aspects of orbit determination, including light-time equations, clock synchronization, station coordinates, and signal propagation delays. For instance, relativistic time delay effects, such as Shapiro delay, can introduce discrepancies of up to tens of microseconds per day, especially in deep-space navigation where signal propagation spans millions of kilometers [9]. Without these corrections, accumulated errors could lead to significant positional offsets, potentially corresponding to meter‐level range errors if uncorrected [9]. A fully consistent treatment across GCRS and BCRS is therefore essential not only to guarantee frame equivalence but also to extend GNSS capabilities to cislunar navigation.
In this work, we report on the implementation of the BCRS framework in JPL’s GipsyX [17], demonstrating millimeter-level closure results. We also derive the necessary coordinate and time transformations needed to define the Lunicentric Celestial Reference System (LCRS) for future Earth–Moon mission applications.
@@ II.1.1 Direct transformations: GCRS $\rightarrow$ BCRS
where $\boldsymbol{\mathrm{x}}_{\tt E}$ and $\boldsymbol{\mathrm{v}}_{\tt E}=d\boldsymbol{\mathrm{x}}_{\tt E}/dt$ being the Earth’s position and velocity vectors in the BCRS, with their magnitudes known to be $x_{\tt E}=|\boldsymbol{\mathrm{x}}_{\tt E}|\equiv{\rm AU}\simeq 1.49\,597\,870\,700\times 10^{11}$ m, and $v_{\tt E}=|\boldsymbol{\mathrm{v}}_{\tt E}|=29\,784.7\,$ m/s, see Table 2. Note that in their most accurate form (see [18]), the time transformations (1) also include terms $\propto{\cal O}(c^{-4})$. The uncertainty in (1) comes from the omitted ${\cal O}(c^{-4})$ terms behaving as $\sim c^{-4}(-\frac{1}{8}v_{\tt E}^{2}-\frac{3}{2}v^{2}_{\tt E}GM_{\tt S}/r_{\tt E}+{\textstyle\frac{1}{2}}(GM_{\tt S}/r_{\tt E})^{2})\lesssim-1.10\times 10^{-16}=-9.50$ ps/d.
@@ II.1.2 Inverse transformations: BCRS $\rightarrow$ GCRS
We can connect the Earth’s center world line, $\boldsymbol{\mathrm{x}}_{\tt E}(t)$, to an observer’s world line, $\boldsymbol{\mathrm{x}}(t)$, using a space-time curve that is orthogonal to both world lines. Note that in data processing programs, the relevant coordinate differences are not simply $\Delta\boldsymbol{\mathrm{r}}$, but rather the differences $\Delta\boldsymbol{\mathrm{r}}=\boldsymbol{\mathrm{x}}(t^{*})-\bar{\boldsymbol{\mathrm{x}}}_{\tt E}(t^{*})$, which are the coordinates at a fixed coordinate time, i.e., simultaneous in solar system barycenter coordinates.
@@ II.2.1 Relativistic time scales
${\tt TT}={\tt TCG}-L_{\tt G}({\tt TCG}-{\tt T}_{0}),\qquad\boldsymbol{\mathrm{X}}_{\tt TT}=(1-L_{\tt G})\boldsymbol{\mathrm{X}}_{\tt TCG},\qquad(GM)_{\tt TT}=(1-L_{\tt G})(GM)_{\tt TCG}.$ (8)
@@ II.3.1 Preliminary derivations
Note that in (18) we also brought in the PPN parameter $\gamma$ as in [1]. We evaluate derivatives along $\tt TT$} unless noted: $\dot{U}_{\rm ext}=-\sum_{\tt B\not=E}GM_{\tt B}\,[(\boldsymbol{\mathrm{v}}_{\tt B}-\boldsymbol{\mathrm{v}}_{\tt E})\cdot(\boldsymbol{\mathrm{x}}_{\tt B}-\boldsymbol{\mathrm{x}}_{\tt E})]/r_{\tt BE}^{3}$ and $\ddot{U}_{\rm ext}$ analogously by differentiating once more; see (27)–(28).
@@ II.4 Practical form for the relevant expressions
As the values for a MEO (GPS‑like) satellite are much larger than those for a ground-based station, we consider the transformations for a MEO orbiter (i.e., a GPS spacecraft), with the position, velocity and the acceleration present in (17)–(18), (33), and (34) take the relevant values of MEO orbiter (i.e., a GPS spacecraft), namely $({\boldsymbol{\mathrm{X}}},\dot{\boldsymbol{\mathrm{X}}},\ddot{\boldsymbol{\mathrm{X}}})\rightarrow({\boldsymbol{\mathrm{X}}}_{\tt MEO},\dot{\boldsymbol{\mathrm{X}}}_{\tt MEO},\ddot{\boldsymbol{\mathrm{X}}}_{\tt MEO})$, that are given in Table 2. With these values, we use (18) to estimate contribution of the various terms in the position transformation:
@@
Equation (36) is referred to as the modified Lorentz transformation. It incorporates the special relativity Lorentz transformation due to Earth’s barycentric velocity, general relativity effects from the gravitational potential, and scaling factors. This transformation affects not only the length of position vectors but also alters the angle between position vectors and the Earth’s velocity vector (aberration). This equation includes more corrective terms than those in ([1], Eq. (4-10)), which JPL’s interplanetary orbit determination software uses for converting station positions from TT-compatible quantities in GCRS to TDB-compatible quantities in BCRS when modeling observables.
@@
Recommended transformations: For cm/ps-class GNSS, use Eq. (36) for position with the two leading $c^{-2}$ terms (scale and Lorentz), Eq. (42) for velocity with the three leading $c^{-2}$ terms, and Eq. (51) for acceleration retaining terms $\gtrsim 10^{-13}\,\mathrm{m\,s^{-2}}$. These screened sets were used to produce the closures in Fig. 1.
@@
matching the remainders from Eqs. (36), (42), and (51) used in our closures. Document these bounds with the same ephemerides to enable external regression.
@@
Implementation note (GM scaling & ephemeris argument): This work adopts TDB as the ephemeris argument. For JPL DE430/DE440-family kernels, the delivered $(GM)$ values are already TDB-consistent; use Eq. (52) only when converting a TT-referenced constant, i.e., $(GM)_{\tt TDB}=(1-L_{\tt C})(GM)_{\tt TT}$ and inversely $(GM)_{\tt TT}=(1-L_{\tt C})^{-1}(GM)_{\tt TDB}$. Do not rescale any $(GM)$ that is tied to a TDB-based kernel.
II.5 Minimal cislunar case study: NRHO-like geometry
@@
Compute one-way light-time between a ground station and the cislunar spacecraft with (64). Retain Sun and Earth Shapiro terms only; other bodies are negligible at the accuracy targeted here. (Note that most of the solar term behaves as a range scale and cancels in differenced observables, but keep it for absolute light-time consistency at the centimeter level over $\Delta T$.) Use the same TT/TDB handling and GM scaling guidance as in Sec. II.4 to avoid double-scaling when using TDB-consistent ephemerides.
@@
and require that the sup norms over $t\in[t_{0},t_{0}+\Delta T]$ do not exceed the screening remainders already specified under Eqs. (36) and (42), respectively. This ensures that all omitted terms remain below the stated operational error budgets when operating in cislunar geometries.
@@ II.6.1 Direct transformations: TDB $\rightarrow$ proper time
where ${\boldsymbol{\mathrm{x}}}_{0}$, $\boldsymbol{\mathrm{v}}_{0}=\dot{\boldsymbol{\mathrm{x}}}_{0}$, and $v_{0}=|\boldsymbol{\mathrm{v}}_{0}|$ represent the BCRS position and velocity vectors of the spacecraft; $\boldsymbol{\mathrm{r}}_{\tt B0}=\boldsymbol{\mathrm{x}}_{B}-\boldsymbol{\mathrm{x}}_{0}$ is the separation vector between a body $B$ and the spacecraft, $r_{\tt B0}=|\boldsymbol{\mathrm{r}}_{\tt B0}|$; also we define $\cos\theta_{\tt B}=(\boldsymbol{\mathrm{n}}_{\tt B0}\cdot\boldsymbol{\mathrm{s}}_{\tt B})$, where $\boldsymbol{\mathrm{s}}_{\tt B}$ is the unit vector along the axis of rotation for the body B. Note that we kept only quadrupole term in $w_{\tt L}$, as contributions of other multipole harmonics of the gravitational fields are typically very small. Ultimately, the error bound in (54) came from the omitted contribution of the $1/c^{4}$–terms (see [19]), that for a GPS spacecraft evaluated to have magnitude of $c^{{}_{4}}(-{\textstyle{1\over 8}}v^{4}-{\textstyle{3\over 2}}v^{2}w_{0}+4(\boldsymbol{\mathrm{v}}\cdot\boldsymbol{\mathrm{w}})+{\textstyle{\textstyle{\textstyle{1\over 2}}}}w_{0}^{2}+\Delta)\simeq 8.77\times 10^{-17}.$ As a result, we may neglect these terms for GNSS applications in cislunar environment.
@@
where $\boldsymbol{\mathrm{r}}_{0}$ and $\boldsymbol{\mathrm{v}}_{0}=d\boldsymbol{\mathrm{r}}_{0}/dt$ are the barycentric position and velocity of the spacecraft, $\boldsymbol{\mathrm{r}}_{\tt B0}=\boldsymbol{\mathrm{r}}_{0}-\boldsymbol{\mathrm{r}}_{\tt B}$ is its position relative to a body $B$, with $\boldsymbol{\mathrm{n}}_{\tt B0}=\boldsymbol{\mathrm{r}}_{\tt B0}/r_{\tt B0}$ being the unit vector in this direction, and $\boldsymbol{\mathrm{s}}_{\tt B}$ is the unit vector alone the axis of the rotation, $\tt TDB_{\tau}$ is the beginning of the integration interval.

### v2

# Relativistic framework for high-precision GNSS processing in GCRS/BCRS with extension to cislunar space

@@ I Introduction
For solar-system navigation beyond Earth orbit it is common to express spacecraft and planetary dynamics in a global barycentric reference system; the IAU-adopted BCRS with coordinate time TCB (or its scaled form TDB) provides the standard realization for modern ephemerides and interplanetary tracking [1, 7, 8]. The mean rate offset between TT and TCB is $L_{\tt B}=1.55051976772\times 10^{-8}$, corresponding to $\simeq 1.34$ ms/day; inconsistent handling of this rate biases modeled one-way and two-way coordinate light-time at the millisecond-per-day level near Earth [9, 8].
As GNSS technology advances with the goal of achieving cm-level positional accuracy, the inclusion of general relativistic corrections in orbit determination models becomes indispensable [10, 11, 12, 13, 14]. Modern cm-class GNSS solutions rely on precise orbit and clock products and precise point positioning (PPP) strategies; see, e.g., [15, 16]. These corrections affect multiple aspects of orbit determination, including light-time equations, clock synchronization, station coordinates, and signal propagation delays [2, 17, 18]. For GNSS links, the dominant Shapiro contribution from the Sun is at the few-nanosecond level (meters in range) with mm-level annual modulation, while Earth’s contribution is at the $\lesssim 50$ ps level (Sec. III.1.3). Inconsistent handling of the TTTCB (or TTTDB) rate introduces a secular $\sim 1.34~{\rm ms/day}$ bias in coordinate light-time near Earth [9, 8]. Without these corrections, accumulated errors could lead to significant positional offsets, potentially corresponding to meter‐level range errors if uncorrected [8].
For interplanetary and lunar applications, navigation observables and dynamical models are often formulated in barycentric coordinates using a first post-Newtonian (11PN) NN-body description consistent with the IAU BCRS metric [19, 1]. The secular rate difference between TT and TCB, $L_{\tt B}-L_{\tt G}=1.5505\times 10^{-8}$, produces a drift of $\simeq 1.34$ ms/day, while the periodic part of ${\tt TDB}-{\tt TT}$ reaches $\simeq 1.6$ ms [9]. These time-scale effects correspond to $\mathcal{O}(10^{2}\!-\!10^{5})\,\mathrm{m}$ one-way range-equivalent biases if left untreated, and thus must be handled explicitly (or absorbed consistently) in any mixed GCRS/BCRS estimation problem. At the link level, the 11PN Shapiro delay contributes up to $\sim 2.7$ ns from the Sun and $\sim 40$ ps from the Earth for a GNSS MEO signal path [5]; near solar conjunction on interplanetary links it can reach $\mathcal{O}(10\!-\!10^{2})\,\mu\mathrm{s}$ [1]. Accordingly, the key requirement for high-precision processing is an internally consistent set of (i) reference-frame and time-scale transformations, (ii) dynamical models, and (iii) signal-propagation models; the particular choice of integration frame (GCRS vs BCRS) is a design choice rather than a physical necessity.
A consistent treatment across GCRS and BCRS is therefore required for cm/ps-class processing when combining geocentric GNSS models with barycentric ephemerides and relativistic light-time. While standard practice propagates GNSS spacecraft in the GCRS and transforms states/observables as needed, a BCRS-native processing option becomes convenient when a single estimation pipeline must also accommodate cislunar or interplanetary geometries.
@@
  1. 1.
    Give explicit closed-form ${\cal O}(c^{-2})$ transformations for position, velocity, and acceleration between TT-compatible GCRS quantities and TDB-compatible BCRS quantities, with conservative remainder bounds and screening thresholds;
  2. 2.
[…]
  3. 3.
[…]
This paper is structured as follows: In Section II, we present the relativistic framework necessary for processing GNSS data from the BCRS, accounting for current measurement precision. This section addresses the relativistic time scales used in astronomical practice (such as TT, TCG, TCB and TDB) and provides expressions to transform positions, velocities, and accelerations between various practically relevant reference frames. In Section III, we demonstrate the practical implementation of processing GNSS data within the BCRS. This section focuses on applying relativistic corrections, including the transformation of observables between reference frames, and addresses the challenges associated with data analysis when accounting for both the BCRS and GCRS systems. Finally, in Section IV, we summarize the results and discuss their implications for future spacecraft navigation, including applications for extending GNSS capabilities beyond Earth’s vicinity and into the Earth-Moon system. In Appendix A, we introduce and discuss the Lunicentric Celestial Reference System (LCRS) with the associated times TCL and TL, which is essential for accurately modeling GNSS receivers operating in cislunar space, where gravitational and relativistic effects must be well-modeled.
@@ II.1.1 Direct transformations: GCRS $\rightarrow$ BCRS
where $\boldsymbol{\mathrm{x}}_{\tt E}$ and $\boldsymbol{\mathrm{v}}_{\tt E}=d\boldsymbol{\mathrm{x}}_{\tt E}/dt$ being the Earth’s position and velocity vectors in the BCRS, with their magnitudes known to be $x_{\tt E}=|\boldsymbol{\mathrm{x}}_{\tt E}|\equiv{\rm AU}\simeq 1.49\,597\,870\,700\times 10^{11}$ m, and $v_{\tt E}=|\boldsymbol{\mathrm{v}}_{\tt E}|=29\,784.7\,$ m/s, see Table 2. Here $t_{0}$ is an arbitrary reference epoch (e.g., J2000) that fixes the integration constant in the time transformation, and $T_{0}\equiv T(t_{0})$ denotes the corresponding TCG epoch. All modeled observables depend only on time differences, so the choice of $t_{0}$ cancels provided it is used consistently throughout the transformation chain. Note that in their most accurate form (see [7]), the time transformations (1) also include terms $\propto{\cal O}(c^{-4})$. The uncertainty in (1) comes from the omitted ${\cal O}(c^{-4})$ terms behaving as $\sim c^{-4}(-\frac{1}{8}v_{\tt E}^{2}-\frac{3}{2}v^{2}_{\tt E}GM_{\tt S}/r_{\tt E}+{\textstyle\frac{1}{2}}(GM_{\tt S}/r_{\tt E})^{2})\lesssim-1.10\times 10^{-16}=-9.50$ ps/d.
@@ II.1.2 Inverse transformations: BCRS $\rightarrow$ GCRS
We can connect the Earth’s center world line, $\boldsymbol{\mathrm{x}}_{\tt E}(t)$, to an observer’s world line, $\boldsymbol{\mathrm{x}}(t)$, using a space-time curve that is orthogonal to both world lines. Note that in data-processing programs the relevant coordinate differences are not simply $\Delta\boldsymbol{\mathrm{r}}$, but rather the simultaneous barycentric differences $\Delta\boldsymbol{\mathrm{r}}=\boldsymbol{\mathrm{x}}(t^{*})-\boldsymbol{\mathrm{x}}_{\tt E}(t^{*})$ evaluated at a common barycentric coordinate time $t^{\ast}$.
@@
Here $T_{0}$ is the conventional reference epoch used in the IAU scaling definitions (see Sec. II.2 and Table 1); the corresponding TCB epoch at the geocenter is $t_{0}=t^{\ast}(T_{0})$.
@@ II.2.1 Relativistic time scales
${\tt TT}={\tt TCG}-L_{\tt G}({\tt TCG}-{\tt T}_{0}),\qquad\boldsymbol{\mathrm{X}}_{\tt TT}=(1-L_{\tt G})\boldsymbol{\mathrm{X}}_{\tt TCG},\qquad(GM)_{\tt TT}=(1-L_{\tt G})(GM)_{\tt TCG},$ (8)
where $T_{0}$ is the conventional reference epoch used in the IAU scaling definitions; in this work we adopt ${\tt T}_{0}=2443144.5003725$ JD, as in Eq. (11).
@@ II.3.1 Preliminary derivations
[…]
@@
Numerically at the geocenter, $GM_{\odot}/{\rm AU}\simeq 8.87\times 10^{8}$ m2/s2 while $GM_{\tt M}/r_{\tt EM}\simeq 1.28\times 10^{4}$ m2/s2; thus the solar term dominates $U_{\rm ext}$ by $\sim 7\times 10^{4}$, with other planetary contributions smaller still.
@@ II.3.2 Transforming velocity and acceleration
The screening and remainder bounds in this section are computed using the parameter envelopes in Table 2, representative of (i) a terrestrial station and (ii) a GNSS MEO spacecraft. For other regimes (e.g., GEO, HEO, lunar transfer, NRHO), the retained terms should be re-screened using the same procedure with updated envelopes for $(\mathbf{X},\dot{\mathbf{X}},\ddot{\mathbf{X}})$ and the relevant external potentials and accelerations.
@@ II.4 Practical form for the relevant expressions
Numerically at the geocenter, $GM_{\tt S}/{\rm AU}\simeq 8.87\times 10^{8}\,{\rm m^{2}/s^{2}}$ while $GM_{\tt M}/r_{\tt EM}\simeq 1.28\times 10^{4}\,{\rm m^{2}/s^{2}}$, so the solar term dominates $U_{\rm ext}$ by $\sim 7\times 10^{4}$ (the remaining planetary terms are smaller still for Earth-orbiting applications).
@@
As the characteristic magnitudes of $(|\boldsymbol{\mathrm{X}}|,|{\dot{\boldsymbol{\mathrm{X}}}}|,|{\ddot{\boldsymbol{\mathrm{X}}}}|)$ for a MEO (GPS-like) spacecraft exceed those of a ground-based station (Table 2), we screen the transformation terms using the MEO envelope to obtain conservative bounds. We consider the transformations for a MEO orbiter (i.e., a GPS spacecraft), with the position, velocity and the acceleration present in (17)–(18), (33), and (34) take the relevant values of MEO orbiter (i.e., a GPS spacecraft), namely $({\boldsymbol{\mathrm{X}}},\dot{\boldsymbol{\mathrm{X}}},\ddot{\boldsymbol{\mathrm{X}}})\rightarrow({\boldsymbol{\mathrm{X}}}_{\tt MEO},\dot{\boldsymbol{\mathrm{X}}}_{\tt MEO},\ddot{\boldsymbol{\mathrm{X}}}_{\tt MEO})$, that are given in Table 2. With these values, we use (18) to estimate contribution of the various terms in the position transformation:
@@
Eq. (36) is closely related to the station-coordinate transformations used in DSN radiometric modeling (e.g., range and Doppler) [1]. Relative to commonly used operational forms, we keep the explicit scale term $L_{\tt C}$ and track the small acceleration-dependent contributions and their remainder bound, which become relevant when demanding mm-level frame-closure consistency and $10^{-16}$-level time/frequency transfer budgets. For GNSS cm-class processing the additional screened terms provide a conservative, implementation-ready state map between TT-compatible GCRS and TDB-compatible BCRS.
@@
[…]
Recommended transformations: For cm/ps-class GNSS, use Eq. (36) for position with the two leading $c^{-2}$ terms (scale and Lorentz), Eq. (42) for velocity with the three leading $c^{-2}$ terms, and Eq. (51) for acceleration retaining terms $\gtrsim 10^{-13}\,\mathrm{m\,s^{-2}}$. These screened sets were used in the 24 h end-to-end frame-closure test of Fig. 1; the bounds below quantify the truncation error introduced solely by screening of the closed-form transformations.
Table 3: Screened operational transformations and conservative remainder bounds using the envelopes in Table 2.
Transformation Position Velocity Acceleration
Screened form ($T\rightarrowTDB$) Eq. (36) Eq. (42) Eq. (51)
Remainder bound (ground) $4.03\times 10^{-6}$ m $9.73\times 10^{-8}$ m/s $5.27\times 10^{-15}$ m/s2
Remainder bound (MEO) $7.01\times 10^{-5}$ m $1.29\times 10^{-7}$ m/s $6.68\times 10^{-14}$ m/s2
@@
matching the remainders from Eqs. (36), (42), and (51) used in our closures. Here $\|\cdot\|$ denotes the Euclidean norm, and $\delta\boldsymbol{\mathrm{r}}(t),\delta\boldsymbol{\mathrm{v}}(t),\delta\boldsymbol{\mathrm{a}}(t)$ are the sums of the omitted terms in the screened position, velocity, and acceleration transformations, respectively. These conservative limits are intended as regression thresholds for the transformations themselves; the end-to-end closure residuals of Sec. III additionally include numerical-integration and force-model consistency errors. For $\Delta T=24$ h, these bounds correspond to worst-case state impacts of $\delta r\sim 7\times 10^{-5}$ m directly from the position remainder, $\delta r\sim(\delta v)\Delta T\approx 1.1$ cm from the velocity remainder, and $\delta r\sim\tfrac{1}{2}(\delta a)\Delta T^{2}\approx 2.5\times 10^{-4}$ m ($0.25$ mm) from the acceleration remainder. Thus the screened acceleration set is consistent with the millimeter-level closure target, while the screened velocity set remains within a centimeter-class envelope over one day.
@@
Implementation note (GM scaling & ephemeris argument): This work adopts TDB as the ephemeris argument. For JPL DE430/DE440-family kernels [33], the delivered $(GM)$ values are already TDB-consistent; use Eq. (52) only when converting a TT-referenced constant, i.e., $(GM)_{\tt TDB}=(1-L_{\tt C})(GM)_{\tt TT}$ and inversely $(GM)_{\tt TT}=(1-L_{\tt C})^{-1}(GM)_{\tt TDB}$. Do not rescale any $(GM)$ that is tied to a TDB-based kernel.
II.5 Minimal cislunar regression test: NRHO-like geometry
@@
Order-of-magnitude check: For an NRHO-like envelope $|\boldsymbol{\mathrm{X}}|\sim 10^{8}$ m and $|\boldsymbol{\mathrm{v}}_{\tt M}|\sim 1$ km/s, the LCRS time-coupling term $c^{-2}(\boldsymbol{\mathrm{v}}_{\tt M}\!\cdot\!\boldsymbol{\mathrm{X}})$ is at the $\sim 10^{-6}$ s level, while the leading spatial correction $\sim c^{-2}v_{\tt M}^{2}\boldsymbol{\mathrm{X}}$ is at the $\sim 10^{-3}$ m level. These scales motivate retaining the full ${\cal O}(c^{-2})$ LCRS terms for cm-level light-time and state modeling in Earth–Moon geometries.
@@
Compute one-way light-time between a ground station and the cislunar spacecraft with (64). Retain Sun and Earth Shapiro terms only; other bodies are negligible at the accuracy targeted here. (Note that most of the solar term behaves as a range scale and cancels in differenced observables, but keep it for absolute light-time consistency at the centimeter level over $\Delta T$.) Use the same TT/TDB handling and GM scaling guidance as in Sec. II.4 to avoid double-scaling when using TDB-consistent ephemerides. The lunar Shapiro delay is negligible at the accuracy targeted here: a rough upper bound is $\Delta t^{\rm Moon}_{\rm Shapiro}\sim 2GM_{\tt M}c^{-3}\ln(\cdots)\lesssim\mathcal{O}(1~{\rm ps})$, corresponding to $\lesssim 0.3$ mm in range, i.e., well below the cm-class absolute light-time envelope adopted in this test.
@@
and require that the sup norms over $t\in[t_{0},t_{0}+\Delta T]$ do not exceed the screening remainders already specified under Eqs. (36) and (42), respectively. In this regression test we adopt the same truncation budgets as for the Earth MEO screening (Table 3) to bound the error introduced solely by the screened state transformations. This ensures that all omitted terms remain below the stated operational error budgets for the specific state and geometry envelope used to size the screened remainder bounds; for mission-specific cislunar operations, the same remainder analysis should be re-evaluated with an NRHO-appropriate envelope for $(|\boldsymbol{\mathrm{X}}|,|\dot{\boldsymbol{\mathrm{X}}}|,|\ddot{\boldsymbol{\mathrm{X}}}|)$.
[…]
@@ II.6.1 Direct transformations: TDB $\rightarrow$ proper time
where ${\boldsymbol{\mathrm{x}}}_{0}$, $\boldsymbol{\mathrm{v}}_{0}=\dot{\boldsymbol{\mathrm{x}}}_{0}$, and $v_{0}=|\boldsymbol{\mathrm{v}}_{0}|$ represent the BCRS position and velocity vectors of the spacecraft; $\boldsymbol{\mathrm{r}}_{\tt B0}=\boldsymbol{\mathrm{x}}_{B}-\boldsymbol{\mathrm{x}}_{0}$ is the separation vector between a body $B$ and the spacecraft, $r_{\tt B0}=|\boldsymbol{\mathrm{r}}_{\tt B0}|$; also we define $\cos\theta_{\tt B}=(\boldsymbol{\mathrm{n}}_{\tt B0}\cdot\boldsymbol{\mathrm{s}}_{\tt B})$, where $\boldsymbol{\mathrm{s}}_{\tt B}$ is the unit vector along the axis of rotation for the body B.
Note that we retained only the quadrupole ($J_{2,{\tt B}}$) contribution in the multipolar correction. Higher harmonics scale as $J_{\ell,{\tt B}}(R_{\tt B}/r_{{\tt B}0})^{\ell}$ and are therefore suppressed by at least $(R_{\tt B}/r_{{\tt B}0})^{2}$ relative to the $J_{2,{\tt B}}$ term. For the Earth at a GPS altitude, $(R_{\tt E}/r)^{2}\simeq 0.058$ and $|J_{4,{\tt E}}/J_{2,{\tt E}}|\sim 10^{-3}$, implying a fractional size $\lesssim 10^{-4}$ of the $J_{2}$ correction, i.e., $\lesssim 10^{-18}$ in $d\tau/dt$, below the $5\times 10^{-18}$ metric cutoff used in Sec. III.2.1. See, e.g., Refs. [5, 7] for conventions and typical harmonic magnitudes. Ultimately, the error bound in (54) came from the omitted contribution of the $1/c^{4}$–terms (see [21]), that for a GPS spacecraft evaluated to have magnitude of $c^{-4}\big(-{\textstyle\frac{1}{8}}v_{0}^{4}-{\textstyle\frac{3}{2}}v_{0}^{2}w_{0}+4(\boldsymbol{\mathrm{v}}_{0}\cdot\boldsymbol{\mathrm{w}})+{\textstyle\frac{1}{2}}w_{0}^{2}+\Delta\big)\simeq 8.77\times 10^{-17}.$ As a result, we may neglect these terms for GNSS applications in cislunar environment.
@@
where $\boldsymbol{\mathrm{r}}_{0}$ and $\boldsymbol{\mathrm{v}}_{0}=d\boldsymbol{\mathrm{r}}_{0}/dt$ are the barycentric position and velocity of the spacecraft, $\boldsymbol{\mathrm{r}}_{\tt B0}=\boldsymbol{\mathrm{r}}_{\tt B}-\boldsymbol{\mathrm{r}}_{0}$ is the separation vector from the spacecraft to body $B$, with $\boldsymbol{\mathrm{n}}_{\tt B0}=\boldsymbol{\mathrm{r}}_{\tt B0}/r_{\tt B0}$ being the corresponding unit vector, and $\boldsymbol{\mathrm{s}}_{\tt B}$ is the unit vector along the axis of the rotation, $\tt TDB_{\tau}$ is the beginning of the integration interval.